A 10-m-long copper wire is pulled with a force of 1300 N and it stretches 10 cm. Calculate the radius of the copper if the value of the Young's modulus is

110 ✕ 109 N/m2.

Assume the cross section of the wire is circular.
Answer #1

We know that young modulus

Y=FL/A∆L

A=FL/Y∆L ..1

here F=1300N , L=10 m , ∆L=10cm=0.10m. Y=110×109 N/m2

put these value in equation 1

A=1300×10/(110×109×0.10)

A=1.182×10-6

since cross section of the wire is circular

πr2=1.182×10-6

r2=0.3762×10

r=0.613×10-3 m

r=0.613 mm .